OpenStudy (anonymous):

There were 37,000 people at a ball game in LA. The day's receipts were $248,000. how many people paid $12. for reserved seats and how many paid $5. for the general admission?

OpenStudy (anonymous):

x= number of general admin tickets y= number of reserve seat tickets

OpenStudy (anonymous):

x+y=37000 5x +12y = 248 000 solve simultaneously

OpenStudy (anonymous):

x+y=37000 5x+12y=248,000

OpenStudy (anonymous):

9,920 paid for general 27,080 paid for reserved

OpenStudy (anonymous):

add them together to get 37,000

OpenStudy (anonymous):

49600 for 5 198,400 for12 added together is 248,000